# Fixed-Time Input-to-State Stability for Singularly Perturbed Systems via Composite Lyapunov Functions

### v3

@@ 1 Introduction
Many complex dynamical systems can be decomposed into components operating on different time scales. Such multi-time-scale systems are common across a diverse array of engineering applications, including aerospace [1], chemical processing [2], smart grids [3], and biological and evolutionary systems [4], among others. Such systems can be studied using techniques from singular perturbation theory [5, 6, 7], where the methods typically focus on decomposing the system into lower order subsystems and studying the subsystems to draw conclusions on the behavior of the overall interconnection.
@@
One simple yet powerful technique for analyzing singularly perturbed systems is the composite Lyapunov method, first introduced by Khalil in [8]. This method is particularly effective for assessing system stability, as it leverages the stability of the subsystems to construct a Lyapunov function for the overall system that is valid under sufficiently large time-scale separation. As demonstrated in [8], if the reduced and boundary layer systems admit certain quadratic-type Lyapunov functions, then under additional interconnection conditions, the interconnected system can be shown to be asymptotically stable, provided there is sufficient time-scale separation between the two subsystems. Moreover, if the Lyapunov functions satisfy specific quadratic bounds, these results can be extended to exponential stability. The composite Lyapunov method has found profound applications in control synthesis across various engineering domains, including power systems [9], biological systems [10], aerospace [11], etc. While the composite Lyapunov method has been widely used to establish asymptotic and exponential stability, it has found limited applications in settings which require stronger notions of stability. In [12], the authors leverage composite Lyapunov functions to establish finite-time stability for a special class of singularly perturbed homogeneous systems. However, to the best of the authors’ knowledge, there does not exist a framework that studies fixed-time stability for a general class of singularly perturbed systems via composite Lyapunov techniques.
@@
Fixed-time stability, which has been popularized through the introduction of Lyapunov conditions in [13], has garnered significant attention due to its ability to address challenges in control [14], optimization [15], and learning [16, 17]. Fixed-time stability ensures convergence to an equilibrium within a fixed time, regardless of the system’s initial conditions. While the Lyapunov conditions simplify the verification of fixed-time stability in continuous-time dynamical systems, techniques for analyzing fixed-time stability in interconnected systems remain limited. Current state-of-the-art methods are primarily applicable to systems with specific homogeneity properties [18], restrictive structural requirements [19], or those satisfying certain small-gain conditions [20]. Moreover, in many practical applications, the systems of interest are also influenced by external disturbances or exogenous inputs. In particular, in this paper we study systems of the form
@@
where $\varepsilon>0$ is a small parameter that induces a time scale separation between the dynamics of $x$ and the dynamics of $z$, and $u$ is an exogenous input. One of the key tools for analyzing such systems is input-to-state stability (ISS), introduced by Sontag in [21] and later characterized through an equivalent Lyapunov framework in [22]. In simple terms, an ISS system converges asymptotically to a bounded set for any bounded input signal. A fixed-time counterpart, known as fixed-time ISS (FxT ISS), was introduced in [23] to address fixed-time stability in systems subject to disturbances. While standard ISS properties have been extensively studied in the context of singularly perturbed systems [24, 25], it remains an open question whether analogous results can be developed for fixed-time stability and whether the composite Lyapunov method introduced in [8] can be extended to analyze FxT ISS.
In this paper, we address the above questions by extending the popular composite Lyapunov method [8] to study FxT ISS in singularly perturbed systems with inputs. Specifically, under the assumptions that the respective “reduced” and “boundary-layer” dynamics of (1) are fixed-time stable on their own, we first derive interconnection conditions under which (1) renders the origin FxT ISS. These interconnection conditions parallel those developed in [8] for asymptotic and exponential stability. However, unlike the results of [8], absence of Lipschitz continuity at the origin results in a more complex paradigm for verifying our proposed interconnection conditions, which involve powers of the fixed-time Lyapunov functions of the reduced and boundary-layer dynamics. Subsequently, we employ our analytical tools to design and analyze a novel class of algorithms that achieve fixed-time feedback optimization under strongly convex cost functions and slowly varying inputs—a class of problems widely studied in the literature [26, 27, 28], but which, to the best of our knowledge, has not previously been addressed using algorithms with fixed-time stability properties. Finally, we present numerical and analytical examples to illustrate the key assumptions and the main stability results of the paper.
@@ 2.2 Auxiliary Results
We first present some lemmas that will be instrumental for our results. The proofs can be found in the Appendix.
@@
where $K:=1+\max\left(1,-\xi_{2}2^{-\xi_{2}-1}\right)$. $\square$
@@ 2.3 Fixed-Time Input-to-State Stability
where $x\in\mathbb{R}^{n}$ is the state and $u\in\mathcal{L}_{\infty}^{p}$ is an input signal. We assume the vector field $f$ is continuous and satisfies $f(0,0)=0$. We will state some definitions from [23] that will be particularly relevant for our work.
@@
It can be verified that the “dissipation” formulation we use in (6) implies relation (10) in [23], which implies FxT ISS for (3) via [23, Thm 4]. Indeed, if (6) holds and we fix some $0<\tilde{\varepsilon}<\min_{i}k_{i}$, we have
@@
where $\chi(\cdot)=\tilde{\chi}^{-1}\circ\rho(\cdot)$, $\tilde{\chi}(\cdot)=\tilde{\varepsilon}\alpha_{2}^{a_{1}}(\cdot)+\tilde{\varepsilon}\alpha_{2}^{a_{2}}(\cdot)$, and $\tilde{k}_{i}=k_{i}-\tilde{\varepsilon}$. By taking $\tilde{\varepsilon}\to 0$, it can be observed via [23, Corollary 2] that systems that admit a FxT ISS Lyapunov function that satisfy (6) also admit a settling time function satisfying the following bound
@@ 3.1 Assumptions
The map $h$ is usually referred to as the quasi-steady state mapping [7] for the disturbance-free system (1b). By using this mapping, we can define the so-called reduced system from (1a):
@@ 3.2 Analysis
where, for simplicity, we omit the arguments of $I_{1}$ and $I_{2}$. With conditions (15a) and (15b), we have:
@@
and we clearly have $\rho_{\varepsilon,\zeta}\in\mathcal{K}_{\infty}$ for $\zeta\in(0,1)$ and $\varepsilon>0$. By (15c) we can find $\zeta^{*}\in(0,1)$ such that $\nu^{*}:=\nu(\zeta^{*})>0$, and then we can find $\varepsilon^{*}>0$ such that $\omega_{\varepsilon}(\zeta^{*})>\nu^{*}$ whenever $\varepsilon\in(0,\varepsilon^{*})$. For these $\varepsilon\in(0,\varepsilon^{*})$ we obtain:
@@
$~~~+\rho_{\varepsilon,\zeta^{*}}(|u|).$
We pick $\gamma_{1}\in\left[\max\left\{a_{1},b_{1}\right\},1\right)$ and $\gamma_{2}\in\left(1,\min\left\{a_{2},b_{2}\right\}\right]$ to obtain
@@
We would like to note that the “or” condition in (15c) is non-exclusive, i.e it is acceptable for both conditions to be satisfied. In this case, any choice of the weight $\zeta\in(0,1)$ will result in a valid composite Lyapunov function candidate.
@@
The functions $\tilde{V}(x)$ and $\tilde{W}(x,y)$ are highly analogous to the functions $\psi_{1}(y)$ and $\psi_{2}(y)$, respectively, from [7, Chapter 11.5]. For the asymptotic stability (resp. FxT ISS) result from [7, Theorem 11.3] (resp. Theorem 1 in this paper), the Lie derivatives of $V$ and $W$ along the reduced and boundary layer dynamics are assumed to be upper bounded by negative multiples of $\psi_{1}^{2}(x)$ and $\psi_{2}^{2}(y)$ (resp. $\tilde{V}^{2}(x)$ and $\tilde{W}^{2}(x,y)$), respectively. Hence, we can observe that the structure of the interconnection conditions in our paper are actually more forgiving than those from [7, Chapter 11.5] in the sense that we allow for extra $\tilde{V}^{2}(x)$ and $\tilde{W}^{2}(x,y)$ terms. This extra degree of freedom is particularly useful since it allows us to apply our results to a variety of interesting systems, such as those presented later in this paper. However, our conditions are also more restrictive in the sense that we now require very specific forms for the expressions $\psi_{1}$ and $\psi_{2}$. But this is expected, since Theorem 1 considers fixed-time ISS, which is a much stronger notion of stability.
We have shown that under suitable assumptions, the inequalities in (15) imply that system (9) is FxT ISS provided there is a sufficiently large time scale separation. The proof of Theorem 1 provides an efficient methodology for computing a somewhat conservative estimate of the required timescale separation, i.e $\varepsilon^{*}$, for FxT ISS. It can also be seen from the proof of Theorem 1 that if the fast dynamics have no input, then the derived gain $\varrho$ in the FxT ISS bound (4) can be made independent of $\varepsilon$. This is further detailed in the following Corollary.
@@
While the results of Theorem 1 and Corollary 2 hold for system (9), the question of whether (1) is FxT ISS may also be of interest. In other words: we ask if the transformation y=zh(x)y=z-h(x) preserves the FxT ISS property. Fortunately, as long as the quasi steady state map satisfies a mild boundedness assumption, (9) being FxT ISS implies (1) is FxT ISS. This is further detailed in the following result.
@@ 3.3 A Stylized Example
where $\lceil\cdot\rfloor^{q}:=|\cdot|^{q}\text{sgn}(\cdot)$, $x,z,u_{1},u_{2}\in\mathbb{R}$, $0<q_{1}\leq r_{1}<1$ and $1<r_{2}\leq q_{2}$. System (21) has the quasi steady state $h(x)=x$, with reduced system:
@@
where $\tilde{q}_{i}=\frac{1}{2}(q_{i}+1)$ and
@@
This implies that system (23) is FxT ISS uniformly in $x$, and hence the assumptions are satisfied. To verify if system (21) is FxT ISS, it remains to check if the interconnection terms associated with system (21) satisfy the interconnection conditions (15). Let $\tilde{V}{(x)}=|x|^{{\tilde{r}_{1}}}+|x|^{{\tilde{r}_{2}}}$ and $\tilde{W}{(y)}=|y|^{{\tilde{q}_{1}}}+|y|^{{\tilde{q}_{2}}}$. Then, we can compute $I_{1}$ and apply Lemmas 1-2 to obtain the following estimate:
$I_{1}$ $=2x\Bigl{(}-\lceil y+x\rfloor^{r_{1}}-\lceil y+x\rfloor^{r_{2}}+\lceil x\rfloor^{r_{1}}+\lceil x\rfloor^{r_{2}}\Bigl{)}$
$\leq c\tilde{V}^{2}{(x)}+2c^{-\sigma_{1}}\tilde{W}^{2}{(y)},$
for all $c\in(0,1)$, where $\sigma_{1}>0$ is obtained by applying Lemma 1 twice and upper bounding appropriately. By taking $c>0$ sufficiently small we can satisfy conditions (15a) and (15c). Thus, it remains to check if $I_{2}$ also satisfies (15b). Indeed, by computing $I_{2}$ we obtain:
@@
$\leq c\tilde{V}^{2}{(x)}+(2c^{-\sigma_{2}}+3+2^{r_{2}})\tilde{W}^{2}{(y)}+|u|^{2},$
for all $c\in(0,1)$, where $\sigma_{2}>0$ is again obtained from Lemma 1. We conclude that the conditions of Theorem 1 and 3 are satisfied, and hence there exists $\varepsilon^{*}$ such that the singularly perturbed system (21) is FxT ISS for $\varepsilon\in(0,\varepsilon^{*})$. We simulate the system using $r_{1}=\frac{2}{5},r_{2}=\frac{6}{5},q_{1}=\frac{1}{3},q_{2}=\frac{9}{7}$ and the disturbances $u_{1}(t)=e^{\sin{t}},u_{2}(t)=\sin(19\log(t+1))-0.21$, where we clearly have $u(t)=[u_{1}(t),u_{2}(t)]^{\top}\in\mathcal{L}_{\infty}^{2}$. The trajectories of this system with and without the input are shown in Figure 1, illustrating the FxT ISS property

### v5

@@ 1 Introduction
Many complex dynamical systems can be decomposed into components operating on different time scales. Such multi-time-scale systems are common across a diverse array of engineering applications, including aerospace [narang2014nonlinear], chemical processing [261484], smart grids [5467283], and biological and evolutionary systems [del2015biomolecular], among others. Such systems can be studied using techniques from singular perturbation theory [vasil1978singular, KokotovicSPBook, khalil], where the methods typically focus on decomposing the system into lower order subsystems and studying the subsystems to draw conclusions on the behavior of the overall interconnection.
@@
One simple yet powerful technique for analyzing singularly perturbed systems is the composite Lyapunov method, first introduced by Khalil in [1103586]. This method is particularly effective for assessing system stability, as it leverages the stability of the subsystems to construct a Lyapunov function for the overall system that is valid under sufficiently large time-scale separation. As demonstrated in [1103586], if the reduced and boundary layer systems admit certain quadratic-type Lyapunov functions, then under additional interconnection conditions, the interconnected system can be shown to be asymptotically stable, provided there is sufficient time-scale separation between the two subsystems. Moreover, if the Lyapunov functions satisfy specific quadratic bounds, these results can be extended to exponential stability. The composite Lyapunov method has found profound applications in control synthesis across various engineering domains, including power systems [9309064], biological systems [9029355], aerospace [naidu2001singular], etc. While the composite Lyapunov method has been widely used to establish asymptotic and exponential stability, it has found limited applications in settings which require stronger notions of stability. For example, in [9992641], the authors leverage composite Lyapunov functions to establish finite-time stability for a special class of singularly perturbed homogeneous systems. However, to the best of the authors’ knowledge, there does not exist a framework that studies fixed-time stability for a general class of singularly perturbed systems via composite Lyapunov techniques.
@@
Fixed-time stability, which has been popularized through the introduction of Lyapunov conditions in [Fixed_timeTAC], has garnered significant attention due to its ability to address challenges in control [polyakov2023finite], optimization [fixed_time], and learning [nonsmoothesc, poveda2022fixed]. Fixed-time stability ensures convergence to an equilibrium within a fixed time, regardless of the system’s initial conditions. While the Lyapunov conditions simplify the verification of fixed-time stability in continuous-time dynamical systems, techniques for analyzing fixed-time stability in interconnected systems remain limited. Current state-of-the-art methods are primarily applicable to systems with specific homogeneity properties [mendoza2023stability], restrictive structural requirements [lei2022event], or those satisfying certain small-gain conditions [10886634]. Moreover, in many practical applications, the systems of interest are also influenced by external disturbances or exogenous inputs. In particular, in this paper we study systems of the form
@@
where $\varepsilon>0$ is a small parameter that induces a time scale separation between the dynamics of $x$ and the dynamics of $z$, and $u$ is an exogenous input. One of the key tools for analyzing such systems is input-to-state stability (ISS), introduced by Sontag in [28018] and later characterized through an equivalent Lyapunov framework in [SONTAG1995351]. In simple terms, an ISS system converges asymptotically to a bounded set for any bounded input signal. A fixed-time counterpart, known as fixed-time ISS (FxT ISS), was introduced in [LOPEZRAMIREZ2020104775] to address fixed-time stability in systems subject to disturbances. While standard ISS properties have been extensively studied in the context of singularly perturbed systems [544001, TeelNesicTAC], it remains an open question whether analogous results can be developed for fixed-time stability and whether the composite Lyapunov method introduced in [1103586] can be extended to analyze FxT ISS.
In this paper, we address the above questions by extending the popular composite Lyapunov method [1103586] to study FxT ISS in singularly perturbed systems with inputs. Specifically, under the assumptions that the respective “reduced” and “boundary-layer” dynamics of (1) are fixed-time stable on their own, we first derive interconnection conditions under which (1) renders the origin FxT ISS. These interconnection conditions parallel those developed in [1103586] for asymptotic and exponential stability. However, unlike the results of [1103586], absence of Lipschitz continuity at the origin results in a more complex paradigm for verifying our proposed interconnection conditions, which involve powers of the fixed-time Lyapunov functions of the reduced and boundary-layer dynamics. Subsequently, we employ our analytical tools to design and analyze a novel class of algorithms that achieve fixed-time feedback optimization under strongly convex cost functions and slowly varying inputs—a class of problems widely studied in the literature [colombino2019online, hauswirth2020timescale, bianchin2022online], but which, to the best of our knowledge, has not previously been addressed using algorithms with fixed-time stability properties. Finally, we present numerical and analytical examples to illustrate the key assumptions and the main stability results of the paper.
@@ 2.2 Auxiliary Results
We first present some lemmas that will be instrumental for our results.
@@
where $K:=1+\max\{1,-\xi_{2}2^{-\xi_{2}-1}\}$. $\square$
@@ 2.3 Fixed-Time Input-to-State Stability
where $x\in\mathbb{R}^{n}$ is the state and $u\in\mathcal{L}_{\infty}^{p}$ is an input signal. We assume the vector field $f$ is continuous and satisfies $f(0,0)=0$. We will state some definitions from [LOPEZRAMIREZ2020104775] that will be particularly relevant for our work.
@@
It can be verified that the “dissipation” formulation we use in (6) implies relation (10) in [LOPEZRAMIREZ2020104775], which implies FxT ISS for (3) via [LOPEZRAMIREZ2020104775, Thm 4]. Indeed, if (6) holds and we fix some $0<\tilde{\varepsilon}<\min_{i}k_{i}$, we have
@@
where $\chi(\cdot)=\tilde{\chi}^{-1}\circ\rho(\cdot)$, $\tilde{\chi}(\cdot)=\tilde{\varepsilon}\alpha_{1}^{a_{1}}(\cdot)+\tilde{\varepsilon}\alpha_{1}^{a_{2}}(\cdot)$, and $\tilde{k}_{i}=k_{i}-\tilde{\varepsilon}$. By taking $\tilde{\varepsilon}\to 0$, it can be observed via [LOPEZRAMIREZ2020104775, Corollary 2] that systems that admit a FxT ISS Lyapunov function that satisfy (6) also admit a settling time function satisfying the following bound
@@ 3.1 Assumptions
The map $h$ is usually referred to as the quasi-steady state mapping [khalil] for the disturbance-free system (1b). By using this mapping, we can define the so-called reduced system from (1a):
@@ 3.2 Analysis
where we used Jensen’s inequality to obtain $2(V^{a_{1}}+V^{a_{2}})\geq\tilde{V}^{2}$ and $2(W^{b_{1}}+W^{b_{2}})\geq\tilde{W}^{2}$, and for simplicity, we omit the arguments of $I_{1}$ and $I_{2}$. With conditions (15a) and (15b), we have:
@@
and we clearly have $\rho_{\varepsilon,\zeta}\in\mathcal{K}_{\infty}$ for $\zeta\in(0,1)$ and $\varepsilon>0$. By (15c) we can find $\zeta^{*}\in(0,1)$ such that $\nu^{*}:=\nu(\zeta^{*})>0$, and then we can find $\varepsilon^{*}>0$ such that $\omega_{\varepsilon}(\zeta^{*})>\nu^{*}$ whenever $\varepsilon\in(0,\varepsilon^{*})$. For these $\varepsilon\in(0,\varepsilon^{*})$, we obtain:
@@
$~~~+\rho_{\varepsilon,\zeta^{*}}(|u|),$
where the second inequality follows by the fact that $(a+b)^{2}\geq a^{2}+b^{2}$ for all $a,b\geq 0$. We pick $\gamma_{1}\in\left[\max\left\{a_{1},b_{1}\right\},1\right)$ and $\gamma_{2}\in\left(1,\min\left\{a_{2},b_{2}\right\}\right]$ to obtain the following inequality via Lemmas 6.1-6.2 in the Appendix:
@@
We would like to note that the “or” condition in (15c) is non-exclusive, i.e., it is acceptable for both conditions to be satisfied. In this case, any choice of the weight $\zeta\in(0,1)$ will result in a valid composite Lyapunov function candidate.
@@
[…]
We have shown that under suitable assumptions, the inequalities in (15) imply that system (9) is FxT ISS provided there is a sufficiently large time scale separation. The proof of Theorem 1 provides an efficient methodology for computing a somewhat conservative estimate of the required timescale separation, i.e., $\varepsilon^{*}$, for FxT ISS. It can also be seen from the proof of Theorem 1 that if the fast dynamics have no input, then the derived gain $\varrho$ in the FxT ISS bound (4) can be made independent of $\varepsilon$. This is further detailed in the following Corollary.
@@
While the results of Theorem 1 and Corollary 2 hold for system (9), the question of whether (1) is FxT ISS may also be of interest. In other words: we ask if the transformation y=zh(x)y=z-h(x) preserves the FxT ISS property. Fortunately, as long as the quasi-steady state map satisfies a mild boundedness assumption, (9) being FxT ISS implies (1) is FxT ISS. This is further detailed in the following result.
@@ 3.3 A Stylized Example
where $\lceil\cdot\rfloor^{q}:=|\cdot|^{q}\text{sgn}(\cdot)$, $x,z,u_{1},u_{2}\in\mathbb{R}$, $0<q_{1}\leq r_{1}<1$ and $1<r_{2}\leq q_{2}$. System (21) has the quasi-steady state $h(x)=x$, with reduced system:
@@
where we used Lemma 6.4 to obtain the first inequality, as well as $2yu_{1}u_{2}\leq c|y|^{2}+\frac{1}{c}|u_{1}u_{2}|^{2}$ (with $c=2^{-q_{2}}$) and $|u_{i}|\leq|u|$ to obtain the second inequality, and
$2^{-q_{2}}|y|^{q_{1}+1}-2^{-q_{2}}|y|^{q_{2}+1}+2^{-q_{2}}|y|^{2}\leq 0$ (25)
to obtain the third inequality with $\tilde{q}_{i}=\frac{1}{2}(q_{i}+1)$ and
@@
This implies that system (23) is FxT ISS uniformly in $x$, and hence the assumptions are satisfied. To verify if system (21) is FxT ISS, it remains to check if the interconnection terms associated with system (21) satisfy the interconnection conditions (15). Let $\tilde{V}{(x)}=|x|^{{\tilde{r}_{1}}}+|x|^{{\tilde{r}_{2}}}$ and $\tilde{W}{(y)}=|y|^{{\tilde{q}_{1}}}+|y|^{{\tilde{q}_{2}}}$. Then, we can compute $I_{1}$:
$I_{1}$ $=2x\Bigl(-\lceil y+x\rfloor^{r_{1}}-\lceil y+x\rfloor^{r_{2}}+\lceil x\rfloor^{r_{1}}+\lceil x\rfloor^{r_{2}}\Bigl).$
By Lemma 2, we have
$\left\lvert\lceil x\rfloor^{r_{1}}-\lceil y+x\rfloor^{r_{1}}\right\rvert\leq 2^{1-r_{1}}|y|^{r_{1}},$ (26)
and
$\left\lvert\lceil x\rfloor^{r_{2}}-\lceil y+x\rfloor^{r_{2}}\right\rvert\leq K|y|\left(|x|^{r_{2}-1}+|y|^{r_{2}-1}\right),$ (27)
where $K=1+\max\{1,(r_{2}-1)2^{r_{2}-2}\}$ and we use the fact that $\lceil x\rfloor^{r}=\frac{x}{|x|^{1-r}}$ for each $r>0$. With (26) and (27), we can use the triangle inequality to bound $I_{1}$ as follows
$I_{1}$ $\leq 2|x|\left(2^{1-r_{1}}|y|^{r_{1}}+K|y|\left(|x|^{r_{2}-1}+|y|^{r_{2}-1}\right)\right)$
$=2^{2-r_{1}}|x||y|^{r_{1}}+2K|y||x|^{r_{2}}+2K|x||y|^{r_{2}}.$ (28)
By Lemma 1, we also have that the following inequalities hold
$2^{2-r_{1}}|x||y|^{r_{1}}$ $\leq 2^{2-r_{1}}\left(\hat{c}|x|^{r_{1}+1}+\hat{c}^{-\frac{1}{r_{1}}}|y|^{r_{1}+1}\right)$ (29a)
$2K|y||x|^{r_{2}}$ $\leq 2K\left(\hat{c}|x|^{r_{2}+1}+\hat{c}^{-r_{2}}|y|^{r_{2}+1}\right)$ (29b)
$2K|x||y|^{r_{2}}$ $\leq 2K\left(\hat{c}|x|^{r_{2}+1}+\hat{c}^{-\frac{1}{r_{2}}}|y|^{r_{2}+1}\right)$ (29c)
for all $\hat{c}>0$. Since $2\tilde{q}_{1}\leq 2\tilde{r}_{1}\leq r_{i}+1\leq 2\tilde{r}_{2}\leq 2\tilde{q}_{2}$ for $i=1,2$, we can use Lemma 6.2 in the Appendix (and the fact that $(a+b)^{2}\geq a^{2}+b^{2},~\forall~a,b\geq 0$) to obtain the following
$|x|^{r_{i}+1}\leq|x|^{2\tilde{r}_{1}}+|x|^{2\tilde{r}_{2}}$ $\leq\tilde{V}^{2}(x)$ (30a)
$|y|^{r_{i}+1}\leq|y|^{2\tilde{q}_{1}}+|y|^{2\tilde{q}_{2}}$ $\leq\tilde{W}^{2}(y)$ (30b)
for $i=1,2$. For fixed $\hat{c}>0$, let $c=(2^{2-r_{1}}+4K)\hat{c}$, and we define the function $\overline{\alpha}:(0,\infty)\to(0,\infty)$ by
$\overline{\alpha}(s)=2^{2-r_{1}}s^{-\frac{1}{r_{1}}}+2K(s^{-r_{2}}+s^{-\frac{1}{r_{2}}}).$
We let $\color[rgb]{0,0,0}\sigma_{1}:=\max\{\frac{1}{r_{1}},r_{2}\}$} and notice that for all $s\in(0,1)$, the function $\overline{\alpha}$ satisfies the following bound
$\overline{\alpha}(s)\leq(2^{2-r_{1}}+4K)s^{-\sigma_{1}}.$ (31)
By combining (28), (29), (30), and (31) we obtain
$I_{1}$ $\leq c\tilde{V}^{2}(x)+\overline{\alpha}(\hat{c})\tilde{W}^{2}(y)$
$=c\tilde{V}^{2}(x)+\overline{\alpha}\left(\frac{c}{2^{2-r_{1}}+4K}\right)\tilde{W}^{2}(y)$
$\leq c\tilde{V}^{2}(x)+{\color[rgb]{0,0,0}(2^{2-r_{1}}+4K)^{1+\sigma_{1}}}c^{-\sigma_{1}}\tilde{W}^{2}(y)$
for all $c\in(0,{\color[rgb]{0,0,0}2^{2-r_{1}}+4K})$. By taking $c>0$ sufficiently small we can satisfy conditions (15a) and (15c). Thus, it remains to check if $I_{2}$ also satisfies (15b). Indeed, by computing $I_{2}$ we obtain:
@@
$\leq 2|y|\left(|y+x|^{r_{1}}+|y+x|^{r_{2}}+|u_{1}|\right)$
$\leq 2|y|\left(|y|^{r_{1}}+|x|^{r_{1}}+2^{r_{2}-1}(|y|^{r_{2}}+|x|^{r_{2}})+|u_{1}|\right)$
$=2(|y|^{r_{1}+1}+|y||x|^{r_{1}}+|y||u_{1}|)+2^{r_{2}}(|y|^{r_{2}+1}+|y||x|^{r_{2}})$
By Lemma 1, we have
$|y||x|^{r_{1}}$ $\leq|x|^{r_{1}+1}+|y|^{r_{1}+1}$
$|y||x|^{r_{2}}$ $\leq|x|^{r_{2}+1}+|y|^{r_{2}+1}.$
Then, we can leverage (30) to obtain
$I_{2}$ $\leq 4|y|^{r_{1}+1}+2|x|^{r_{1}+1}+|y|^{2}+|u|^{2}+2^{r_{2}+1}|y|^{r_{2}+1}$
$~~~+2^{r_{2}}|x|^{r_{2}+1}$
$\leq{\color[rgb]{0,0,0}(2+2^{r_{2}})}\tilde{V}^{2}(x)+{\color[rgb]{0,0,0}(5+2^{r_{2}+1})}\tilde{W}^{2}(y)+|u|^{2}.$
We conclude that the conditions of Theorem 1 and 3 are satisfied, and hence there exists $\varepsilon^{*}$ such that the singularly perturbed system (21) is FxT ISS for $\varepsilon\in(0,\varepsilon^{*})$. We simulate the system using $r_{1}=\frac{2}{5},r_{2}=\frac{6}{5},q_{1}=\frac{1}{3},q_{2}=\frac{9}{7}$ and the disturbances $u_{1}(t)=e^{\sin{t}},u_{2}(t)=\sin(19\log(t+1))-0.21$, where we clearly have $u(t)=[u_{1}(t),u_{2}(t)]^{\top}\in\mathcal{L}_{\infty}^{2}$. The trajectories of this system with and without the input are shown in Figure 1, illustrating the FxT ISS property